# An Exploratory Study on Human-Robot Interaction using Semantics-based Situational Awareness

### v1

@@ IV-B Correlation analysis
We employ Spearman’s correlation analysis to find the hidden relationship/patterns among these factors. The correlation heat map of each measurement is presented in Fig. 7.
With the comparison of the two heatmaps, we observe some interesting findings with strong correlations. They are: a) The parallel task accuracy has a positive mild correlation with the NASA TLX ($\rho$ = 0.40) in the semantics trial, i.e. when the parallel task accuracy is high, the cognitive workload is high. However, this shows a negative correlation ($\rho$ = -0.21) in the non-semantics condition. I.e. when the parallel task accuracy is high, the cognitive workload appears slightly decreased. b) The parallel task completion speed is significantly negatively correlated with parallel task accuracy degradation ($\rho$ = -0.71) in the semantics trial. This suggests that, when the parallel task completion speed is high, the parallel task accuracy is higher than baseline. However, this becomes non-correlated ($\rho$ = -0.06) in the non-semantics trial. c) The trust in the system has a mild negative correlation with the NASA TLX ($\rho$ = -0.37) in the semantics trial. I.e. when trust in the system is high, cognitive workload is low. However, this shows non-correlated in the non-semantics trial ($\rho$ = 0.05). d) Trust in the system has a mild negative correlation with the LoA switch number ($\rho$ = -0.43) in the non-semantics trial. This suggests that, when trust in the system is high, the number of LoA switches is low. However, this shows little correlation ($\rho$ = 0.03) in the semantics trial.
@@ V Discussion
In summary, the significance analysis suggests that using high-level semantic indicators, based on our framework can significantly benefit HRT with respect to: response time to switch LoA; workload of the human operator; and the operator’s trust in their awareness and understanding of the robot’s environment. High-level semantic indicators also had a modest impact on the number of victims found and the human’s trust in the robotic system.
@@
  • a) The correlation between parallel task accuracy and the NASA TLX suggests that the high-level semantic indicators may be providing useful additional information, but this requires additional cognitive workload. If participants seek to improve the parallel task performance by exploiting the semantic indicator displays, this can cause additional workload. Hence, even though high-level semantics helps to alleviate cognitive workload, this may cause the human to exert additional effort to improve the parallel task accuracy, beyond what might otherwise have been possible without the semantic indicators.
  • b) Correlation between trust in the system and NASA TLX (with-semantics trial) suggests that, when participants have a lower workload, they may trust the system more. However, the non-semantics trial showed no such correlation. Perhaps this suggests that participants who felt a cognitive load reduction when using semantic indicators, felt a strong sense of trust in the indicators element of the system specifically.
  • c) Trust shows negative correlation with number of LoA switches (non-semantic trials), but little correlation in with-semantics trials. Participants with low system trust make many LoA switches (perhaps “noisy decisionmaking”[29]) when semantic indicators are unavailable. However, introducing semantic indicators seems to have a smoothing effect on this behaviour, regardless of the participant’s statements about their own sense of trust.
The box plots of Fig. 6 suggest a correlation between high trust in the robotic system, and large amounts of driving the robot in teleoperation mode. Meanwhile item c) above suggests an interesting relationship between trust and LoA switching, prompting us to explore these relationships further. From Fig. 8 we see that participants with low trust and no semantics strongly favour autonomy mode. In contrast, with semantic indicators, partipants with low trust strongly favour teleoperation. This suggests that: i) the semantic indicator displays comprise a large part of participants’ SA, i.e. displaying these semantic indicators causes participants to feel they have a much better understanding of the remote scene surrounding the robot; ii) when participants trust in their SA, they feel much more confident about using teleoperation mode; iii) they prefer teleoperation mode (provided they have sufficient SA) when they have low trust in the robotic “system” i.e. in the autonomous agent. Furthermore, we can see that, at high levels of trust, participants converge on equal use of autonomy and teleoperation, regardless of the amount of SA available to them.
@@ VI Limitations
To minimize learning effects as a confounder in our “within-groups” experimental design, we used (slightly) different arena layouts between the two trials of each participant. This might have introduced a difference in overall difficulty between the two trials, although we tried to design the two arenas to have similar difficulties. To check this, we asked participants which arena they thought was more difficult. Five participants thought both arenas were equally easy. Three thought the with-semantics arena was more difficult. Six thought the without-semantics arena was more difficult. While not identical, these results suggest similar difficulties. Note also that the without-semantics condition might make the without-semantics arena seem more difficult, further reassuring that the arenas had similar overall difficulty.
Several participants felt that their performance would improve with a wider-angle camera, a finding supported in the literature [24]. Participants with no experience in video games may struggle to understand the 2D bird’s-eye view (map) and easily become lost in the arena, despite a substantial pre-experiment training session.
Some participants waited for the sound warning from the system (3 seconds) to end and before they switched LoA. This unnecessarily increased the human idle time in the experiment. Other methods, such as short vibrations on the joystick, might be a better option for delivering this semantic warning.
@@ VII Conclusion
In this paper, we validated the usefulness of our proposed semantic indicators framework [9], and investigated the effects of high-level semantic indicators in HRT and HRI. Moreover, we explored variable autonomy HRT patterns when high-level semantics are involved. Our experiments indicate that displaying high-level semantic indicators can help humans: decrease reaction time when switching LoA, reduce cognitive workload, and increase trust in their situational awareness.

### v2

@@ IV-B Correlation analysis
We employ Spearman’s correlation analysis to find the hidden relationship/patterns among these factors. The correlation heat map of each measurement is presented in Fig. 7. With the comparison of the two heatmaps, we observe some interesting findings with strong correlations. They are:
  • The parallel task accuracy has a positive mild correlation with the NASA TLX ($\rho$ = 0.40) in the semantics trial, i.e. when the parallel task accuracy is high, the cognitive workload is high. However, this shows a negative correlation ($\rho$ = -0.21) in the non-semantics condition. I.e., when the parallel task accuracy is high, the cognitive workload appears slightly decreased.
  • The parallel task completion speed is significantly negatively correlated with parallel task accuracy degradation ($\rho$ = -0.71) in the semantics trial. This suggests that, when the parallel task completion speed is high, the parallel task accuracy is higher than baseline. However, this becomes non-correlated ($\rho$ = -0.06) in the non-semantics trial.
  • The trust in the system has a mild negative correlation with the NASA TLX ($\rho$ = -0.37) in the semantics trial. I.e. when trust in the system is high, cognitive workload is low. However, this shows non-correlated in the non-semantics trial ($\rho$ = 0.05).
  • Trust in the system has a mild negative correlation with the LoA switch number ($\rho$ = -0.43) in the non-semantics trial. This suggests that, when trust in the system is high, the number of LoA switches is low. However, this shows little correlation ($\rho$ = 0.03) in the semantics trial.
@@ V Discussion
The significance analysis indicates that high-level semantics have impacts on response time to switch LoA, NASA TLX, and situational trust scale. Specifically, with the help of the semantic indicators, human operators could quickly switch LoA ($W$ = 16.0, $P$ = 0.04, $r$ = 0.570) when the situation changes (e.g. new personal belongings detected). And, NASA TLX significantly decreases ($W$ = 11.0, $P$ = 0.007, $r$ = 0.724), indicating that semantic indicators help to reduce overall cognitive load. Moreover, the situational trust scale significantly increases ($W$ = 11.5, $P$ = 0.031, $r$ = 0.577), which means human operators have more trust in their SA when assisted by semantic indicators.
Besides the above significant changes, the analysis shows the trend that high-level semantics might have positive effects on the number of victims found ($W$ = 0, $P$ = 0.102, $r$ = 0.436) and trust in the system ($W$ = 10.5, $P$ = 0.080, $r$ = 0.468). These trends have the potential to become significant impacts if we involve more participants in the experiment. In the end, no significant differences were detected in rest indicators (parallel task accuracy, completion rate, navigation completion, etc.), suggesting that simply adding high-level semantics does not improve these performances. They may still be affected by task design and other factors in the experiment.
In summary, the significance analysis suggests that using high-level semantic indicators, based on our framework can significantly benefit HRT with respect to: response time to switch LoA; workload of the human; and the operator’s trust in their SA. High-level semantic indicators shows a trend that they help to increase the number of victims found and human’s trust in the robotic system.
@@
  • The correlation between parallel task accuracy and the NASA TLX suggests that the high-level semantic indicators may be providing useful additional information, but this requires additional cognitive workload. If participants seek to improve the parallel task performance by exploiting the semantic indicator displays, this might cause additional workload ($\rho$ = 0.40 suggests a moderate positive monotonic relationship). Hence, even though high-level semantics helps alleviate cognitive workload, this may cause the human to exert additional effort to improve the parallel task accuracy, beyond what might otherwise have been possible without the semantic indicators. A similar finding is claimed by [24] that “subjects may have been able to distribute their excess attention to other displays and tasks”, if the automation improves SA by reducing excessive workload.
  • Correlation between trust in the system and NASA TLX (with-semantics trial) suggests that, when participants have a lower workload, they may trust the system more. However, the non-semantics trial shows no such correlation. By providing the high-level semantics, human participants could strengthen their trust in the system from the saved “bandwidth” of cognitive load [25].
  • Trust shows negative correlation with the number of LoA switches (non-semantic trials), but little correlation in with-semantics trials. Some researchers [26] claim that “higher trust usually results in an operator accepting the automation’s decisions.” Human operators take more control when trust goes down. However, we believe that there are more complex reasons hidden behind. Participants with low system trust make many LoA switches (perhaps “noisy decision making” [27]) when semantic indicators are unavailable. Noisy decision making could easily happen when humans are in a high-uncertainty situation (poor SA). By introducing semantic indicators (lowering uncertainty and improving SA), human operators might avoid dropping into that high-uncertainty situation.
The box plots of Fig. 6 suggest a correlation between high trust in the robotic system, and large amounts of driving the robot in teleoperation mode. Meanwhile, the correlation between trust and the number of LoA switches in different experiments suggests an interesting relationship, prompting us to explore these relationships further. In Fig. 8, we see that participants with low trust and no semantics strongly favour autonomy mode. In contrast, with semantic indicators, participants with low trust strongly favour teleoperation. This suggests that: i) the semantic indicator displays comprise a large part of participants’ SA, i.e. displaying these semantic indicators causes participants to feel they have a much better understanding of the remote scene surrounding the robot; ii) when participants trust in their SA, they feel much more confident about using teleoperation mode; iii) they prefer teleoperation mode (provided they have sufficient SA) when they have low trust in the robotic “system” i.e. in the autonomous agent. Furthermore, we can see that, at high levels of trust, participants converge on equal use of autonomy and teleoperation, regardless of the amount of SA available to them.
@@ VI Limitations
To minimize learning effects as a confounder in our “within-groups” experimental design, we used (slightly) different arena layouts between the two trials. This might have introduced a difference in overall difficulty between the two trials, although we tried to design the two arenas to have similar difficulties. To check this, we asked participants which arena they thought was more difficult. Five participants thought both arenas were equally easy. Three thought the with-semantics arena was more difficult. Six thought the without-semantics arena was more difficult. While not identical, these results suggest similar difficulties. Note also that the without-semantics condition might make the without-semantics arena seem more difficult, further reassuring that the arenas had similar overall difficulty.
Some participants waited for the sound warning (3 seconds) to end and before they switched LoA. This unnecessarily increased the human idle time. It indicates multimodal notification, such as vibrations on the joystick, could be a better option for delivering this semantics.
@@ VII Conclusion
In this paper, we validated the usefulness of our proposed framework [9], and investigated the effects of high-level semantic indicators in HRT and HRI. Moreover, we explored VA HRT patterns when high-level semantics are involved. Our experiments indicate that displaying high-level semantic indicators can help humans: decrease reaction time when switching LoA, reduce cognitive workload, and increase trust in their SA.